# Soft Condorcet Optimization for Ranking of General Agents

### v2

@@ C.3 Relationship to Perturbed Optimizers
In particular, consider the case where $\mathcal{C}$ is a convex polytope and the problems are linear programs. As an example, the problem of ranking a vector of $d$ entries can be cast in this fashion, by taking $C$ to be the permutahedron $\mathcal{P}(\rho)$ (see below). The function $\mathbf{y}^{*}:\mathbb{R}^{d}\to\mathcal{Y}$ is piecewise constant, leading to zero or undefined gradients. It can be smoothed by taking the mean under additive stochastic noise to the parameters: we define this proxy by the perturbed maximizer (Berthet et al., 2020) for $\varepsilon>0$
@@
where $Z$ is a random variable with a positive and differentiable density. Under these conditions, the argmax inside the expectation is almost surely unique, and $\mathbf{y}^{*}_{\varepsilon}$ is differentiable, with derivatives that are also expressed as a mean, and that can be estimated with Monte-Carlo methods (see (Berthet et al., 2020)). Further, one can define the probability distribution $p_{\bm{\theta}}(y)=\Pr(\mathbf{y}^{*}(\bm{\theta}+\varepsilon Z)=y)$. It holds that $\mathbf{y}^{*}_{\varepsilon}(\bm{\theta})=\mathbb{E}_{p_{\bm{\theta}}(y)}[Y]$, and the perturbed maximum is
@@
Ranking can be cast as an LP in the form required by equation (24). Let $\mathcal{P}(\mathbf{w})$ be the permutahedron: the convex hull over all vectors that are permutations of $\mathbf{w}\subset\mathbb{R}^{m}$. For $\rho=(m,\cdots,1)$, a (linear) objective function is constructed that is the dot product of $-\bm{\theta}$ and $\mathbf{y}\in\mathcal{P}(\rho)$. The fundamental theorem of linear programming ensures that the solution of an LP is vertex of it polytope, unique for almost all inputs, corresponding to a permutation in the case of a permutahedron (Berthet et al., 2020). As an example, suppose the ratings for agents $1,2,3,4$ are $(10,20,30,40)$. Then the maximum objective would be when $\mathbf{y}=\rho$ with a maximal value of $(4,3,2,1)\cdot(-10,-20,-30,-40)$. These methods learn such ranking functions from data via gradient descent. Indeed $\mathbf{y}_{\varepsilon}^{*}(\bm{\theta})$ can be plugged in directly into any loss $\ell$ comparing it to a vector of ranks $\mathbf{y}_{\text{true}}$. Further, the formalism of Fenchel-Young losses (Blondel et al., 2020a) can be applied to this setting: they are particularly convenient to use with perturbed optimizer, as their gradient is given by
@@
Appendix D Additional Results and Hyper-Parameter Values
D.1 Full SCO Results on the Failure Mode of Elo
@@
Table 4: Number of iterations to reach convergence to the optimal ranking on the example in Equation 10. $\alpha$ is the learning rate and $\tau$ is the temperature of the logistic function (Equation (5)). $T_{conv}$ is the number of iterations required to converge to the optimal ranking; for SGD this is an average over 3 seeds.
D.2 Kemeny-Young Approximation Quality
@@
D.3 Approximation of Bayesian Posterior
Figure 5: Posterior inference via Algorithm 1 on a synthetic evaluation problem. Contests between $4$ participants are randomly drawn from a pool of $10$. The ranking of the final rating $\theta_{T}$ is denoted with an asterisk (*). The true skill ranking is labeled in bold. True skills ratings, to two decimal places, are [126.46,106.00,114.68,133.61,128.01,85.34,114.25,97.73,98.45,[126.46,106.00,114.68,133.61,128.01,85.34,114.25,97.73,98.45,[ 126.46 , 106.00 , 114.68 , 133.61 , 128.01 , 85.34 , 114.25 , 97.73 , 98.45 , 106.16]106.16]106.16 ]. All rankings sampled from the SGD-as-posterior-inference process are labeled in the bar plot with their corresponding observed frequencies as probabilities displayed above them. Notice that the mode of the posterior matches the true skill ranking and improves upon the final iterate.
@@ D.4 Online Performance
Figure 6: Online performance of SCO. Bands around each line represent 95% confidence intervals over 50 runs.

### v3

@@ D.3. Relationship to Perturbed Optimizers
In particular, consider the case where $\mathcal{C}$ is a convex polytope and the problems are linear programs. As an example, the problem of ranking a vector of $d$ entries can be cast in this fashion, by taking $C$ to be the permutahedron $\mathcal{P}(\rho)$ (see below). The function $\mathbf{y}^{*}:\mathbb{R}^{d}\to\mathcal{Y}$ is piecewise constant, leading to zero or undefined gradients. It can be smoothed by taking the mean under additive stochastic noise to the parameters: we define this proxy by the perturbed maximizer Berthet et al. (2020) for $\varepsilon>0$
@@
where $Z$ is a random variable with a positive and differentiable density. Under these conditions, the argmax inside the expectation is almost surely unique, and $\mathbf{y}^{*}_{\varepsilon}$ is differentiable, with derivatives that are also expressed as a mean, and that can be estimated with Monte-Carlo methods (see Berthet et al. (2020)). Further, one can define the probability distribution $p_{\bm{\theta}}(y)=\Pr(\mathbf{y}^{*}(\bm{\theta}+\varepsilon Z)=y)$. It holds that $\mathbf{y}^{*}_{\varepsilon}(\bm{\theta})=\mathbb{E}_{p_{\bm{\theta}}(y)}[Y]$, and the perturbed maximum is
@@
Ranking can be cast as an LP in the form required by equation (25). Let $\mathcal{P}(\mathbf{w})$ be the permutahedron: the convex hull over all vectors that are permutations of $\mathbf{w}\subset\mathbb{R}^{m}$. For $\rho=(m,\cdots,1)$, a (linear) objective function is constructed that is the dot product of $-\bm{\theta}$ and $\mathbf{y}\in\mathcal{P}(\rho)$. The fundamental theorem of linear programming ensures that the solution of an LP is vertex of it polytope, unique for almost all inputs, corresponding to a permutation in the case of a permutahedron Berthet et al. (2020). As an example, suppose the ratings for agents $1,2,3,4$ are $(10,20,30,40)$. Then the maximum objective would be when $\mathbf{y}=\rho$ with a maximal value of $(4,3,2,1)\cdot(-10,-20,-30,-40)$. These methods learn such ranking functions from data via gradient descent. Indeed $\mathbf{y}_{\varepsilon}^{*}(\bm{\theta})$ can be plugged in directly into any loss $\ell$ comparing it to a vector of ranks $\mathbf{y}_{\text{true}}$. Further, the formalism of Fenchel-Young losses Blondel et al. (2020a) can be applied to this setting: they are particularly convenient to use with perturbed optimizer, as their gradient is given by
@@
Appendix E Additional Results and Hyper-Parameter Values
E.1. Full SCO Results on the Failure Mode of Elo
@@
Table 3. Number of iterations to reach convergence to the optimal ranking on the example in Equation 11. $\alpha$ is the learning rate and $\tau$ is the temperature of the logistic function (Equation (5)). $T_{conv}$ is the number of iterations required to converge to the optimal ranking; for SGD this is an average over 3 seeds.
E.2. Kemeny-Young Approximation Quality
@@
E.3. Approximation of Bayesian Posterior
Figure 5. Posterior inference via Algorithm 1 on a synthetic evaluation problem. Contests between $4$ participants are randomly drawn from a pool of $10$. The ranking of the final rating $\theta_{T}$ is denoted with an asterisk (*). The true skill ranking is labeled in bold. True skills ratings, to two decimal places, are [126.46,106.00,114.68,133.61,128.01,85.34,114.25,97.73,98.45,[126.46,106.00,114.68,133.61,128.01,85.34,114.25,97.73,98.45,[ 126.46 , 106.00 , 114.68 , 133.61 , 128.01 , 85.34 , 114.25 , 97.73 , 98.45 , 106.16]106.16]106.16 ]. All rankings sampled from the SGD-as-posterior-inference process are labeled in the bar plot with their corresponding observed frequencies as probabilities displayed above them. Notice that the mode of the posterior matches the true skill ranking and improves upon the final iterate.
@@ E.4. Online Performance
Figure 6. Online performance of SCO. Bands around each line represent 95% confidence intervals over 50 runs.